# Self-generated electrokinetic flows from active-charged boundary patterns

### v1

@@ III.4 Long wavelength approximation
The streamfunction satisfies the boundary conditions (no-slip velocity field)
@@
Here we look at the ionic flux and surface charge variation patterns in Eqs (1) and (2), respectively, with a large (but finite) wavelength compared to the Debye length, like $l\sim 10^{2}\lambda-10^{5}\lambda$ for instance. Keeping terms up to order $\mathcal{O}(\tilde{k}^{-1})$, we get the analytical expressions for the leading order streamfunction (see Appendix A)
@@
where $\Psi_{0}(\tilde{r})$ and $\Psi_{1}(\tilde{r})$ are radially-dependent functions whose explicit form is displayed in Appendix B. It is clear that the streamfunction (28) displays two distinct behaviors. The first term on the right hand-side gives rise to a unidirectional velocity component (a zero mode), which is a natural consequence of the respective body force (first term in the square brackets of Eq. (20)), cf. panel (c) of Fig. 2. The remaining two terms in the streamfunction (28) give rise to circulatory behavior, in accordance with the last two terms in the square brackets of Eq. (20), cf. panel (b) of Fig. 2. Cross term effects have appeared before in the literature [22], for a two-dimensional channel, albeit with respect to a formulation different than ours. The electric body force in their case was linearized with respect to a basic state (of uniform surface charge and ionic flux conditions) and thus gave rise to circulating flows only without any zero mode. In contrast, here we establish the existence of a zero mode velocity component and circulatory flows, the latter as a consequence of wall flux activity alone. The resulting flow field and net velocity from our stream function are discussed in the next sections, where we analyze the selection rules revealing their specific dependencies on key parameters.
IV Flow states generated by selective flux-charge phase differences
Active-charged surfaces inside the capillary produce a variety of flow states with a rich dependence on the flux and charge phase differences resulting in various selection rules for $\alpha_{\pm}$ and $\beta$. Using the stream function Eq. (28) derived above, we find the leading order radial and axial fluid velocity components from Eq. (24) as
@@ IV Flow states generated by selective flux-charge phase differences
and $\Psi_{0}(\tilde{r})$ and $\Psi_{1}(\tilde{r})$ were defined in Appendix B. The factors $U(\tilde{r})$ and $W(\tilde{r})$ are odd with respect to $\tilde{k}$. Thus, under spatial inversion the velocity $\tilde{v}_{z}$ changes sign. The form of Eq. (30) was predicted in Eq. (3) by symmetry principles.
@@
A basic condition that leads to a steady flow state, even with vanishing surface charge $\sigma_{0}=0$ ($\delta_{\sigma}=0$), is when only one of the ionic flux, say the cationic flux is active $j_{0+}\neq 0$ ($\delta_{+}\neq 0$) while the anionic flux is set to zero $j_{0-}=0$ ($\delta_{-}=0$). Under this condition, only the circulatory part remains as $h_{s}\neq 0$, but the unidirectional part vanishes since $h_{0}=0$. In the converse situation with only the anionic flux active and centered at origin instead, leads to the exact same circulatory states. Another case with no surface charge that also results in such circulatory flows is when both flux are active $j_{0\pm}\neq 0$ given that their relative amplitudes are not the same which means $\delta_{+}\neq\delta_{-}$. But if $\delta_{+}=\delta_{-}$ then the cationic-anionic flux phase difference needs to be $\alpha_{-}\neq n\pi/2$ (for some integer $n$) such that it gives non-vanishing circulating flows. However, in all these scenarios with vanishing surface charge, only circulatory states can arise. We will now look what happens when a flux-charge phase difference is introduced with non-vanishing surface charge pattern $\sigma_{0}\neq 0$ ($\delta_{\sigma}\neq 0$).
IV.1 Illustrative example: flow with only cationic activity
@@
Lets first look at the effect of just the phase difference $\beta$ between cationic flux and surface charge by setting the anionic flux to zero $j_{0-}=0$, that is $\delta_{-}=0$ in Eqs. (21)-(23). Here if we also set flux-charge amplitudes such that $\delta_{+}=\delta_{\sigma}$, the coefficients of the flow velocity turn out as $h_{0}\sim\sin(\beta)$, $h_{c}\sim-\sin(\beta)$, and $h_{s}\sim 1-\cos(\beta)$. Then when the flux and charge patterns are in-phase (same phase) $\beta=2n\pi$, all the coefficients vanish leading no flow as $v_{r,z}\approx 0$ implying that the fluid becomes stagnant. But if they are out-of-phase (opposite phase) $\beta=(2n+1)\pi$, the unidirectional component vanishes and the axial velocity becomes $v_{z}\sim W\sin(kz)$. Hence in this case, we only develop fully circulatory flow states of toroidal vortex patterns that alternate in direction with profile shapes of ±$\pm W$ across the capillary (see Fig. 2(b)). However, if the flux-charge patterns are offset by an arbitrary phase difference that is not of the same or opposite phase, then a unidirectional contribution emerges. For instance, when $\beta=\pi/2$ the unidirectional contribution is maximal as $v_{z}\sim U+W[\sin(kz)-\cos(kz)]$ leading to a unidirectional dominant flow diminishing the vortex patterns (see Fig. 2(c)). But if the relative contributions of unidirectional and circulatory components are comparable, like when $\beta=3\pi/4$, it results in intermediate states with alternating expanded and contracted vortex patterns where along the capillary it has alternating profile shapes $U\pm W$ (see Fig. 2(d)).
@@
Figure 3: Self-generated electrokinetic flows from flux-charge patterns with both cationic and anionic activity. The stream plots of fluid velocity in $(r,z)$ space obtained analytically from Eq. (29) and (30) and numerically using COMSOL (Appendix C), show good agreement. The modulation-pattern configurations employ $\alpha_{+}=0$ and $\alpha_{-}=\pi/2$ at different flux-charge phase difference, leading to (a) unidirectional dominant flow at $\beta=0$, (b) fully circulatory flow at $\beta=3\pi/4$ and (c) intermediate flow state at $\beta=\pi/2$. The corresponding modulation forms of the cationic (+) flux, the anionic (-) flux and the surface charge are shown above. The arrows of streamlines indicate the flow direction, and the color map indicates the velocity magnitude. (d) Averaged capillary flow velocity plot with the wavelength averaged z-component velocity as a function of flux-charge phase difference $\beta$ with $\alpha_{+}=0$ and $\alpha_{-}$ as indicated, comparing the analytical given by Eq. (33) with the numerical computed in COMSOL (Appendix C). All the quantities are given in terms of the Debye length $\lambda$ for bulk concentration $C_{\infty}=1$ mM. We use here capillary radius $R=10\lambda$ and wavelength $l=200\lambda$ for ions with same diffusivity $D_{+}=D_{-}$ and flux-charge peak amplitudes such that $\delta_{+}=\delta_{-}=\delta_{\sigma}=10^{-4}$.
@@
This effect arising from the flux-charge phase difference can be physically understood as spatial mismatch or defects appearing in the surface structure that constantly injects electrochemical gradients and triggers fluid flow as a response. For instance, here with the anionic flux set to zero, the parameter $\beta$ represents the degree of mismatch between the sites of peak cation in-flux (out-flux) and peak positive (negative) surface charge. When in-phase like $\beta=0$, the peak sites are in perfect alignment and the patterns fully overlap such that the dislocated charges balances out. As a result, body force inducing electrochemical gradients vanish as fluid settles to rest. In the case of an anti-alignment of the peak sites like for $\beta=\pi$, that is cation in-flux peak matches with negative surface charge peak and vice versa. This type of opposing overlap incurs a constant redistribution of charge and ends up in a steady circulating flow. However if offset with an arbitrary misalignment of the flux-charge peaks like $\beta=\pi/2$, breaks spatial symmetry of translation-reflection and leads to transverse body forces that persist, invoking steady flows with a preferred net direction. Moreover, in a biological context this can be envisioned as a scenario of elongated cell membranes with an alternating sequence of gated ion pumps that facilitate in-and out-flux of cations, and a lipid bilayer surface that carries heterogeneous distribution of charge. This can also be realized in a laboratory construction by periodic coatings of ion-absorbing and ion-releasing chemical reactants grafted on micro-channels or capillary walls in addition with alternating induction of surface charges.
@@
Additionally we note that in the converse case of only anionic activity ($j_{0-}\neq 0$) with cationic flux now set to zero ($j_{0+}=0$), the flow directions are reversed and states are changed according to $\beta$ with the peak anionic flux placed at the center. Specifically, in Eqs. (21)-(23) for $\delta_{+}=0$ and $\alpha_{-}=0$, we obtain the two coefficients with flipped signs $h_{0,c}\rightarrow-h_{0,c}$, whereas the other one goes as $h_{s}\rightarrow 2-h_{s}$. This means that when the ionic activity is swapped (from cationic only to anionic only), for a flux-charge phase difference like $\beta=\pi/2$ of the former case (Fig. 2(c)), the unidirectional and the circulatory parts in the latter case turn to the opposite directions reversing the flow. However, for the intermediate state seen previously at $\beta=3\pi/4$ (Fig. 2(c)), the corresponding reversed scenario with similar vortex structure would now be that of $\beta=\pi/4$ and not $\beta=3\pi/4$ for this case here. The swapping of ionic activity can also leads to a shift in flow states. Furthermore, the consequence of being in-phase or out-of-phase are switched here, namely for same phase $\beta=2n\pi$ there is only circulatory flows and instead for opposite phase $\beta=(2n+1)\pi$ there are no flows.
IV.2 Illustrative example: flow with both cationic and anionic activity
@@
Next we look at the scenario where both fluxes are non-zero and examine the resulting flow patterns induced by the phase differences. Figures 3(a)-(c) illustrates the fluid velocity when ion diffusivity and flux amplitudes are identical, $D_{+}=D_{-}$ and $j_{0+}=j_{0-}$ such that the relative flux and charge amplitudes are all the same $\delta_{+}=\delta_{-}=\delta_{\sigma}$. In this case if the cationic and anionic fluxes are in-phase $\alpha_{-}=2\pi n$, then there is no unidirectional component as $h_{0}=0$ in Eq. (21) regardless of what the flux-charge phase difference $\beta$ is. Thus leading to only circulatory flows producing the toroidal vortex patterns. However, if the $\pm$ fluxes are offset with some arbitrary $\alpha_{-}$, we get a non-vanishing unidirectional component depending on $\beta$ and results in a variety of composite flow states. For instance when $\alpha_{-}=\pi/2$, the unidirectional part dominates at $\beta=0$ diluting the circulatory vortex patterns (Fig. 3(a)), whereas in contrast, unidirectionality subsides at $\beta=3\pi/4$ giving fully circulatory flows with amplified vortex patterns (Fig. 3(b)). We observe the alternating vortex patterns at intermediary flow states here for $\beta=\pi/2$ (Fig. 3(c)). In addition, we validate our analytical results with numerical finite element simulations as shown in Figs. 3(a)-(c) using the COMSOL Multiphysics software (see Appendix C for details).
@@
With both ionic fluxes being non-zero, the phase difference $\alpha_{-}$ can be thought of as an extent of spatial separation between two types of active sites, one that produces cationic flux and the other produces anionic flux. For the special case when the peak $\pm$ fluxes are fully aligned, like at $\alpha_{-}=0$, this can be taken as scenario of a surface activity that simultaneously produces fluxes of both ionic species at the same site. If the relative cationic-anionic flux amplitudes are also the same, then this can only result in circulating flows independent of how the surface charge pattern is offset. Here the effect of the cationic flux-charge phase difference $\beta$ discussed previously (in Sec. IVA) that gives rise to the unidirectional flow gets counteracted by the exact opposite effect from anionic flux-charge phase difference which in this case is also $\beta$. Thus the effect of flux-charge phase difference is suppressed when the cites of cationic-anionic activity (in- and out-flux) fully coincide. However for some arbitrary spatial discrepancies in the peak cationic-anionic active sites such as with $\alpha_{-}=\pi/2$ that is not in-phase, the opposing effects are now mismatched with some non-vanishing net unidirectional component. The degree of this mismatch can then be tuned by the offset of the charge pattern, resulting again in a spectrum of flow states.
@@
Taken together, a spectrum of steady states from circulatory, unidirectional and intermediate flows are obtained through active-charged modulation patterns on confinement boundaries. Flow states with a non-vanishing unidirectional component emerge when ionic flux and surface charge patterns are offset by a flux-charge phase difference attributed to a spatial mismatch in the peak sites of flux and charge distributions. This can be achieved with just one ionic (cationic or anionic) activity (Sec. IVA) as well as for the case with both (Sec. IVB). In the former case, the spectrum of flow states are produce when the flux and charge patterns mismatched by a phase difference such that they are neither of the same or opposite phase $\beta\neq n\pi$. Moreover, swapping out the type of ionic activity here (from cationic to anionic, or vice versa) leads to a complete flow reversal or a change of the flow patterns for a fixed phase difference. In the latter case, the simultaneous presences of equal cationic and anionic flux at the same cite suppresses unidirectional flow, and an offset between the two fluxes $\alpha_{-}\neq 2n\pi$ is required to recover the full spectrum of flow states. We focused in this section on the effect of phase difference parameters by setting the relative flux and charge amplitudes as equivalent. We will explore in the Sec. VI the effect of have different relative flux magnitudes by the means of ion diffusivity mismatch. In the following section we characterize the averaged fluid flow across the capillary and provide estimates for the net unidirectional velocity obtained as a function of the key parameters.
@@
Debye length (bulk ion concentration) Amplitudes of flux, charge {$j_{0+}$ (mM m sec-1), $\sigma_{0}$ (C m-2)} Relative amplitudes (dimensionless) $\delta_{+,\sigma}$ Average (peak) flow velocity for capillary radius $R\sim 10^{-1}$ $\mu$m , wavelength $l\sim 1-10^{2}$ $\mu$m
@@
$\lambda\sim 10$ nm ($C_{\infty}=1$ mM) {$10^{-4}$, $10^{-6}$} $10^{-3}$ $\langle v_{z}\rangle\sim 10^{-1}-10$ $\mu$m sec-1
{$10^{-3}$, $10^{-5}$} $10^{-2}$ $\langle v_{z}\rangle\sim 10-10^{3}$ $\mu$m sec-1
{$10^{-2}$, $10^{-4}$} $10^{-1}$ $\langle v_{z}\rangle\sim 10^{3}-10^{5}$ $\mu$m sec-1
@@
$\lambda\sim 1$ nm ($C_{\infty}=10^{2}$ mM) {$10^{-1}$, $10^{-5}$} $10^{-3}$ $\langle v_{z}\rangle\sim 1-10^{2}$ $\mu$m sec-1
{$1$, $10^{-4}$} $10^{-2}$ $\langle v_{z}\rangle\sim 10^{2}-10^{4}$ $\mu$m sec-1
{$10$, $10^{-3}$} $10^{-1}$ $\langle v_{z}\rangle\sim 10^{4}-10^{6}$ $\mu$m sec-1
@@
Table 1: Estimates of the average unidirectional (peak) velocity given by Eq. (33), with only cationic flux active ($j_{0-}=0$) for radius $R=0.1$ $\mu$m and wavelength in the range of $l\sim 1-100$ $\mu$m, evaluated at different Debye lengths $\lambda$ (for corresponding bulk ion concentrations $C_{\infty}$) and peak amplitudes of cationic flux ($j_{0+}$ in units of m sec-1) and surface charge ($\sigma_{0}$ in units of m-2). The values of the flux and charge amplitudes taken here are all small, that is the (dimensionless) amplitudes in Eq. (18) are $\delta_{+,\sigma}\sim 10^{-3}$, $10^{-2}$, and $10^{-1}$.

### v2

@@ III.4 Long wave mode
The streamfunction satisfies the boundary conditions (no-slip and no-penetration velocity field)
@@
Here we look at the ionic flux and surface charge variation patterns in Eqs (1) and (2), respectively, with a large (but finite) wavelength compared to the Debye length, like $l\sim 0.1-10^{2}$ $\mu$m for instance. Keeping terms up to order $\mathcal{O}(\tilde{k}^{-1})$, we get the analytical expressions for the leading order streamfunction (see Appendix A)
@@
where $\Psi_{0}(\tilde{r})$ and $\Psi_{1}(\tilde{r})$ are radially-dependent functions whose explicit form is displayed in Appendix C. It is clear that the streamfunction (28) displays two distinct behaviors. The first term on the right hand-side gives rise to a unidirectional velocity component (a zero mode), which is a natural consequence of the respective body force (first term in the square brackets of Eq. (20)), cf. panel (c) of Fig. 2. The remaining two terms in the streamfunction (28) give rise to circulatory behavior, in accordance with the last two terms in the square brackets of Eq. (20), cf. panel (b) of Fig. 2. Cross term effects have appeared before in the literature [23], for a two-dimensional channel, albeit with respect to a formulation different than ours. The electric body force in their case was linearized with respect to a basic state (of uniform surface charge and ionic flux conditions) and thus gave rise to circulating flows only without any zero mode. In contrast, here we establish the existence of a zero mode velocity component and circulatory flows, the latter as a consequence of wall flux activity alone. The resulting flow field and net velocity from our stream function are discussed in the next sections, where we analyze the selection rules revealing their specific dependencies on key parameters.
IV Flux-charge phase difference
Active-charged patterns on the capillary wall produce a variety of flow states with a rich dependence on the flux and charge phase differences resulting in various selection rules for $\alpha_{\pm}$ and $\beta$. Using the stream function Eq. (28) derived above, we find the leading order radial and axial fluid velocity components from Eq. (24) as
@@ IV Flux-charge phase difference
and $\Psi_{0}(\tilde{r})$ and $\Psi_{1}(\tilde{r})$ were defined in Appendix C. The factors $U(\tilde{r})$ and $W(\tilde{r})$ are odd with respect to $\tilde{k}$. Thus, under spatial inversion the velocity $\tilde{v}_{z}$ changes sign. The form of Eq. (30) was predicted in Eq. (3) by symmetry principles.
@@
A basic condition that leads to a steady flow state, even with vanishing surface charge $\sigma_{0}=0$ ($\delta_{\sigma}=0$), is when only one of the ionic flux, say the cationic flux is active $j_{0+}\neq 0$ ($\delta_{+}\neq 0$) while the anionic flux is set to zero $j_{0-}=0$ ($\delta_{-}=0$). Under this condition, only the circulatory part remains as $h_{s}\neq 0$, but the unidirectional part vanishes since $h_{0}=0$. In the converse situation with only the anionic flux active and centered at origin instead, leads to the exact same circulatory states. Another case with no surface charge that also results in such circulatory flows is when both fluxes are active $j_{0\pm}\neq 0$ given that their relative amplitudes are not the same which means $\delta_{+}\neq\delta_{-}$. But if $\delta_{+}=\delta_{-}$ then the cationic-anionic flux phase difference needs to be $\alpha_{-}\neq n\pi/2$ (for some integer $n$) such that it gives non-vanishing circulating flows. However, in all these scenarios with vanishing surface charge, only circulatory states can arise. We will now look what happens when a flux-charge phase difference is introduced with non-vanishing surface charge pattern $\sigma_{0}\neq 0$ ($\delta_{\sigma}\neq 0$).
IV.1 Case with only cationic activity
@@
Lets first look at the effect of just the phase difference $\beta$ between cationic flux and surface charge by setting the anionic flux to zero $j_{0-}=0$, that is $\delta_{-}=0$ in Eqs. (21)-(23). Here if we also set flux-charge amplitudes such that $\delta_{+}=\delta_{\sigma}$, the coefficients of the flow velocity turn out as $h_{0}\sim\sin(\beta)$, $h_{c}\sim-\sin(\beta)$, and $h_{s}\sim 1-\cos(\beta)$. Then when the flux and charge patterns are in-phase (same phase) $\beta=2n\pi$, all the coefficients vanish leading no flow as $v_{r,z}\approx 0$ implying that the fluid becomes stagnant. But if they are out-of-phase (opposite phase) $\beta=(2n+1)\pi$, the unidirectional component vanishes and the axial velocity becomes $\tilde{v}_{z}\sim W\sin(\tilde{k}\tilde{z})$. Hence in this case, we only develop fully circulatory flow states of toroidal vortex patterns that alternate in direction with profile shapes of ±$\pm W$ across the capillary (see Fig. 2(b)). However, if the flux-charge patterns are offset by an arbitrary phase difference that is not of the same or opposite phase, then a unidirectional contribution emerges. For instance, when $\beta=\pi/2$ the unidirectional contribution is maximal as $\tilde{v}_{z}\sim U+W[\sin(\tilde{k}\tilde{z})-\cos(\tilde{k}\tilde{z})]$ leading to a unidirectional dominant flow diminishing the vortex patterns (see Fig. 2(c)). But if the relative contributions of unidirectional and circulatory components are comparable, like when $\beta=3\pi/4$, it results in intermediate states with alternating expanded and contracted vortex patterns where along the capillary it has alternating profile shapes $U\pm W$ (see Fig. 2(d)).
@@
Figure 3: Self-generated electrokinetic flows from flux-charge patterns with both cationic and anionic activity. The stream plots of fluid velocity in $(\tilde{r},\tilde{z})$ space obtained analytically from Eq. (29) and (30) and numerically using COMSOL (Appendix D), show good agreement. The modulation pattern configurations employ $\alpha_{+}=0$ and $\alpha_{-}=\pi/2$ at different flux-charge phase difference, leading to (a) unidirectional dominant flow at $\beta=0$, (b) fully circulatory flow at $\beta=3\pi/4$ and (c) intermediate flow state at $\beta=\pi/2$. The corresponding modulation forms of the cationic (+) flux, the anionic (-) flux and the surface charge are shown above. The arrows of streamlines indicate the flow direction, and the color map indicates the velocity magnitude where $\tilde{r}=r/\lambda$, $\tilde{z}=z/\lambda$ and $\bar{v}=v/\lambda$ s-1. (d) Averaged capillary flow velocity plot with the wavelength averaged z-component velocity as a function of flux-charge phase difference $\beta$ where $\langle\bar{v}_{z}\rangle=\langle v_{z}\rangle/\lambda$ s-1 with $\alpha_{+}=0$ and $\alpha_{-}$ as indicated, comparing the analytical result given by Eq. (33) with the numerical result computed in COMSOL (Appendix D). The quantities are given in terms of the Debye length $\lambda$ for bulk concentration $C_{\infty}=1$ mM. We use here capillary radius $R=10\lambda$ and wavelength $l=200\lambda$ with flux-charge peak amplitudes such that $\delta_{\pm}\approx\delta_{\sigma}\approx 3.2\times 10^{-4}$ for ions with same diffusivity $D_{\pm}=10^{-9}$ m2s-1.
@@
This effect arising from the flux-charge phase difference can be physically understood as spatial mismatch or defects appearing in the surface structure that constantly injects electrochemical gradients and triggers fluid flow as a response. For instance, here with the anionic flux set to zero, the parameter $\beta$ represents the degree of mismatch between the sites of peak cation in-flux (out-flux) and peak positive (negative) surface charge. When in-phase like $\beta=0$, the peak sites are in perfect alignment and the patterns fully overlap such that the dislocated charges balances out. As a result, body force inducing electrochemical gradients vanish as fluid settles to rest. In the case of an anti-alignment of the peak sites like for $\beta=\pi$, that is cation in-flux peak matches with negative surface charge peak and vice versa. This type of opposing overlap incurs a constant redistribution of charge and ends up in a steady circulating flow. However an offset with an arbitrary misalignment of the flux-charge peaks like $\beta=\pi/2$, breaks spatial symmetry of translation-reflection and leads to transverse body forces that persist, invoking steady flows with a preferred net direction. Moreover, in a biological context this can be envisioned as a scenario of elongated cell membranes with an alternating sequence of gated ion pumps that facilitate in-and out-flux of cations, and a lipid bilayer surface that carries heterogeneous distribution of charge. This can also be realized in a laboratory construction by periodic coatings of ion-absorbing and ion-releasing chemical reactants grafted on micro-channels or capillary walls in addition with alternating induction of surface charges.
@@
Additionally we note that in the converse case of only anionic activity ($j_{0-}\neq 0$) with cationic flux now set to zero ($j_{0+}=0$), the flow directions are reversed and states are changed according to $\beta$ with the peak anionic flux placed at the center. Specifically, in Eqs. (21)-(23) for $\delta_{+}=0$ and $\alpha_{-}=0$, we obtain the two coefficients with flipped signs $h_{0,c}\rightarrow-h_{0,c}$, whereas the other one goes as $h_{s}\rightarrow 2-h_{s}$. This means that when the ionic activity is swapped (from cationic only to anionic only), for a flux-charge phase difference like $\beta=\pi/2$ of the former case (Fig. 2(c)), the unidirectional and the circulatory parts in the latter case turn to the opposite directions reversing the flow. However, for the intermediate state seen previously at $\beta=3\pi/4$ (Fig. 2(c)), the corresponding reversed scenario with similar vortex structure would now be that of $\beta=\pi/4$ and not $\beta=3\pi/4$ here. The swapping of ionic activity can also leads to a shift in flow states. Furthermore, the consequence of being in-phase or out-of-phase are switched here, namely for same phase $\beta=2n\pi$ there is only circulatory flows and instead for opposite phase $\beta=(2n+1)\pi$ there are no flows.
IV.2 Case with both cationic and anionic activity
@@
Next we look at the scenario where both fluxes are non-zero and examine the resulting flow patterns induced by the phase differences. Figures 3(a)-(c) illustrates the fluid velocity when ion diffusivity and flux amplitudes are identical, $D_{+}=D_{-}$ and $j_{0+}=j_{0-}$ such that the relative flux and charge amplitudes are all the same $\delta_{+}=\delta_{-}=\delta_{\sigma}$. In this case if the cationic and anionic fluxes are in-phase $\alpha_{-}=2\pi n$, then there is no unidirectional component as $h_{0}=0$ in Eq. (21) regardless of what the flux-charge phase difference $\beta$ is. Thus leading to only circulatory flows producing the toroidal vortex patterns. However, if the $\pm$ fluxes are offset with some arbitrary $\alpha_{-}$, we get a non-vanishing unidirectional component depending on $\beta$ and results in a variety of composite flow states. For instance when $\alpha_{-}=\pi/2$, the unidirectional part dominates at $\beta=0$ diluting the circulatory vortex patterns (Fig. 3(a)), whereas in contrast, unidirectionality subsides at $\beta=3\pi/4$ giving fully circulatory flows with amplified vortex patterns (Fig. 3(b)). We observe the alternating vortex patterns at intermediary flow states here for $\beta=\pi/2$ (Fig. 3(c)). In addition, we validate our analytical results with numerical finite element simulations as shown in Figs. 3(a)-(c) using the COMSOL Multiphysics software (see Appendix D for details).
@@
With both ionic fluxes being non-zero, the phase difference $\alpha_{-}$ can be thought of as an extent of spatial separation between two types of active sites, one that produces cationic flux and the other produces anionic flux. For the special case when the peak $\pm$ fluxes are fully aligned, like at $\alpha_{-}=0$, this can be taken as scenario of a surface activity that simultaneously produces fluxes of both ionic species at the same site. If the relative cationic-anionic flux amplitudes are also the same, then this can only result in circulating flows independently of how the surface charge pattern is offset. Here the effect of the cationic flux-charge phase difference $\beta$ discussed previously (in Sec. IVA) that gives rise to the unidirectional flow gets counteracted by the exact opposite effect from anionic flux-charge phase difference which in this case is also $\beta$. Thus the effect of flux-charge phase difference is suppressed when the cites of cationic-anionic activity (in- and out-flux) fully coincide. However for some arbitrary spatial discrepancies in the peak cationic-anionic active sites such as with $\alpha_{-}=\pi/2$ that is not in-phase, the opposing effects are now mismatched with some non-vanishing net unidirectional component. The degree of this mismatch can then be tuned by the offset of the charge pattern, resulting again in a spectrum of flow states.
@@
Taken together, a spectrum of steady states from circulatory, unidirectional and intermediate flows are obtained through active-charged modulation patterns on confinement boundaries. Flow states with a non-vanishing unidirectional component emerge when ionic flux and surface charge patterns are offset by a flux-charge phase difference attributed to a spatial mismatch in the peak sites of flux and charge distributions. This can be achieved with just one ionic (cationic or anionic) activity (Sec. IVA) as well as for the case with both (Sec. IVB). In the former case, the full spectrum of flow states are produce when the flux and charge patterns mismatched by a phase difference such that they are neither of the same or opposite phase $\beta\neq n\pi$. Moreover, swapping out the type of ionic activity here (from cationic to anionic, or vice versa) leads to a complete flow reversal or a change of the flow patterns for a fixed phase difference. In the latter case, the simultaneous presences of equal cationic and anionic flux at the same cite suppresses unidirectional flow, and an offset between the two fluxes $\alpha_{-}\neq 2n\pi$ is required to recover the full spectrum of flow states. We focused in this section on the effect of phase difference parameters by setting the relative flux and charge amplitudes as equivalent. We will explore in the Sec. VI the effect of having different relative flux magnitudes by the means of ion diffusivity mismatch. In the following section we characterize the averaged fluid flow across the capillary and provide estimates for the net unidirectional velocity obtained as a function of the key parameters.
@@
Debye length (bulk ion concentration) Flux amplitude $j_{0+}$ (mM m s-1) Charge amplitude $\sigma_{0}$ (C m-2) Relative amplitudes (dimensionless) for $\delta_{+}\approx\delta_{\sigma}$ Average (peak) flow velocity for radius $R=0.5$ $\mu$m and wavelength $l=1$ $\mu$m
@@
$\lambda\approx 14$ nm ($C_{\infty}=1$ mM) $2.3\times 10^{-4}$ $4.2\times 10^{-6}$ $\delta_{+,\sigma}\approx 3.2\times 10^{-3}$ $\langle v_{z}\rangle\approx 0.1$ $\mu$m s-1
$7.3\times 10^{-3}$ $1.3\times 10^{-4}$ $\delta_{+,\sigma}\approx 0.1$ $\langle v_{z}\rangle\approx 10^{3}$ $\mu$m s-1
@@
$\lambda\approx 1.4$ nm ($C_{\infty}=10^{2}$ mM) $2.3\times 10^{-1}$ $4.2\times 10^{-5}$ $\delta_{+,\sigma}\approx 3.2\times 10^{-3}$ $\langle v_{z}\rangle\approx 1$ $\mu$m s-1
$7.3$ $1.3\times 10^{-3}$ $\delta_{+,\sigma}\approx 0.1$ $\langle v_{z}\rangle\approx 10^{4}$ $\mu$m s-1
@@
Table 1: Estimates of the average unidirectional (peak) velocity given by Eq. (33), with only cationic flux active for radius $R=0.5$ $\mu$m and wavelength of $l=1$ $\mu$m, evaluated at different Debye lengths $\lambda$ (for corresponding bulk ion concentrations $C_{\infty}$) and peak amplitudes of cationic flux ($j_{0+}$ in units of m s-1) and surface charge ($\sigma_{0}$ in units of m-2). The values of the flux and charge amplitudes taken here are small, that is the (dimensionless) amplitudes in Eq. (18) are $\delta_{+}\approx\delta_{\sigma}\lesssim 0.1$ and $kR\approx 3$.